OpenStudy (anonymous):

complete the square: x^2-4x+2=0 i need detailed steps showing how to solve.. im lost on this one

OpenStudy (saifoo.khan):

x^2-4x+2 = 0 x^2 - 4x = -2 x^2 - 4x + (4/2)^2 - (4/2)^2 = -2 (x - 2)^2 - (2)^2 = -2 (x-2)^2 - 4 = -2 (x - 2)^2 -4 + 2 = 0 (x-2)^2 - 2 = 0